OpenStudy (patrickweewoos):

Carissa and Louann both have savings accounts. Carissa has $250 in her account and deposits $80 per month. Louann has $1230 in her account and withdraws $60 per month. After how many months will they have equal amounts of money in their accounts? What amount of money will each have when it is the same amount? It will take_______ months. Then, they will both have $_______ in their accounts.

OpenStudy (patrickweewoos):

@perl and @iGreen

OpenStudy (anonymous):

lol I just asked this same exact question...

OpenStudy (perl):

Rasheedah did you tag ?

OpenStudy (patrickweewoos):

Really lol

OpenStudy (perl):

gotta tag multiple people

OpenStudy (igreen):

You can write two equations, graph them and then find out the solution. I answered this question yesterday..lol.

OpenStudy (patrickweewoos):

LOL srry @perl

OpenStudy (anonymous):

no i didnt tag

OpenStudy (igreen):

Carrisa: 250 + 80x Louann: 1230 - 60x Graph these two expressions here: https://www.desmos.com/calculator

OpenStudy (perl):

C = 250 + 80x L = 1230 - 60x

OpenStudy (igreen):

Just copy and paste them into separate boxes on the left. @PatrickWeeWoos

OpenStudy (perl):

solve C = L

OpenStudy (igreen):

Or that^

OpenStudy (perl):

@iGreen what do you mean ?

OpenStudy (perl):

you can find old posts?

OpenStudy (igreen):

Yeah..if you google the questions..but I answered this question yesterday.

OpenStudy (patrickweewoos):

Solve C =L huh?

OpenStudy (perl):

solve C =L which means solve 250 + 80x = 1230 - 60x

OpenStudy (perl):

x = 7

OpenStudy (patrickweewoos):

Tht does not =1230

OpenStudy (perl):

It will take 7 months, and they both have 810 in their account

OpenStudy (patrickweewoos):

@perl

OpenStudy (perl):

It will take___7____ months. Then, they will both have $___810____ in their accounts.

OpenStudy (perl):

close question please

OpenStudy (perl):

i fanned and testimonied you

OpenStudy (patrickweewoos):

Thx @perl

OpenStudy (secret-ninja):

@RasheedahOwl I was helping you in your question. :[